Thanks for the answers for the previous questions. Pls, am having difficulty with these ones. I need your help. it's very urgent


4x^3 - 5x. is it even, odd, or neither?


-4x^2 - 5. is it odd, even or neither?


f(x) = -4x divided by |5x|. is it odd, even, or neither?


-2x^3 divided by 7x^2 + 8. is it odd, even, or neither?


f(x)= 8x^3 + 3. is it odd, even, or neither?

Answer by stanbon(75887)

4x^3 - 5x. is it even, odd, or neither?


-4x^2 - 5. is it odd, even or neither?
-f(-x) = -[-4x^2-5] = 4x^2+5 so not odd
-----------------------------
f(-x) = -4x^2-5 so even
-----------------------------

f(x) = -4x divided by] |5x|. is it odd, even, or neither?
f(x) = -f(-x) so odd
-----------------------------


-2x^3 divided by 7x^2 + 8. is it odd, even, or neither?
-f(-x) = -2(-x)^3 / 7(-x)^2+8 = 2x^3/(7x^2+8)
not odd; not even; so neither
---------------------------------


f(x)= 8x^3 + 3. is it odd, even, or neither?
-f(-x) = -[-8x^3+3] = = 8x^3-3
not odd; not even; so neither.
=================================
Cheers,
Stan H.